The outboard motor of a banca has an efficiency of 22% and it discards 1 x 10³] of heat every second. How much work does the motor do every second?

1
Expert's answer
2022-02-10T15:12:45-0500

Heat discarded every second = 1000 which represent 88% what about 100%?

heat energy supplied "=\frac{100}{88} \times 1000=1136.3636"

efficiency"=\frac{work \space done}{heat \space energy \space supplied} \times 100\%"

"22=\frac{work\space done} {1136.3636}\times 100"

work done "=1136.3636\times \frac{22}{100}=250J"